• retagged by
3,693 views
26 26 votes

An electric circuit between two terminals $A$ and $B$ is shown in the figure below, where the numbers indicate the probabilities of failure for the various links, which are all independent.

What is the probability that $A$ and $B$ are connected?

  1. $\left(\dfrac{6}{25}\right)$
  2. $\left(\dfrac{379}{400}\right)$
  3. $\left(\dfrac{1}{1200}\right)$
  4. $\left(\dfrac{1199}{1200}\right)$
  5. $\left(\dfrac{59}{60}\right)$

3 Answers

Best answer
28 28 votes

Lets define the following events:

$P$ : Uppermost link is working.

$Q$: Middle link is working.

$R$: Lowermost link is working.

$W$ : Terminals $A$ & $B$ are connected.


From the given information, we can calculate the probabilities of events $P$, $Q$ and $R$ as follow:

(as I already consumed letter (capital) $P$ denoting an event so here by I will use letter (small) $p$ to denote probabilities,

e.g. $p\left ( X \right )$ denotes probability of happening of event $X$).

$\qquad p\left ( P \right ) = \dfrac{4}{5} \times \dfrac{4}{5} = \dfrac{16}{25}$,

$\Rightarrow\; p\left ( \widehat{P} \right ) = 1 - \dfrac{16}{25} = \dfrac{9}{25}$,

$\qquad p\left ( Q \right ) = \dfrac{2}{3} $,

$\Rightarrow\; p\left ( \widehat{Q} \right ) = \dfrac{1}{3} $,

$\qquad p\left ( R \right ) = \dfrac{3}{4} \times \dfrac{3}{4} = \dfrac{9}{16}$,

$\Rightarrow\; p\left ( \widehat{R} \right ) = 1 - \dfrac{9}{16} = \dfrac{7}{16}$.


The situation here can be represented using the following Venn Diagram:

Here Red region denotes the event $\widehat{W}$ in which $A \& B$ are disconnected.

Different gradients of Green colour represent the $W$ in which terminals $A \& B$ are connected.

We have to find $p\left ( W \right )$

Events $W$ and $\widehat{W}$ are mutually exclusive & totally exhaustive,

since either terminals $A$ and $B$ will be connected or they will be disconnected.

So we can write $p\left ( W \right )$ in terms of $p\left ( \widehat{W} \right )$ as follow:

$ p\left ( W \right )= 1 - p\left ( \widehat{W} \right )$.

Also it can be seen that 

$ p\left ( \widehat{W} \right )= p\left ( \widehat{P} \right ) \cap \left ( \widehat{Q} \right ) \cap \left ( \widehat{R} \right )$ where $\widehat{P}$, $\widehat{Q}$ and $\widehat{R}$ are independent events.

That is terminals $A$ & $B$ will be disconnected only when all of the links will fail simultaneously.

Using independence we can write,

$p\left ( \widehat{W} \right ) = p\left ( W \right )= p\left ( \widehat{P} \right ) \times \left ( \widehat{Q} \right ) \times \left ( \widehat{R} \right )$,

$\Rightarrow p\left ( \widehat{W} \right )= \dfrac{9}{25} \times \dfrac{1}{3} \times \dfrac{7}{16} = \dfrac{21}{400}$

$\Rightarrow p\left ( W \right )= 1 -  \dfrac{21}{400} = \dfrac{379}{400}$.

Correct Answer: $B$

• edited by
19 19 votes

If component are connected in Series(Both working) then we can say Probability of A connected to B
P(A ∩ B) = P(A)P(B)
If component are connected in Parallel(any one among two working) then we can say Probability of A connected to B
P(A U B) = P(A)+P(B) - P(A∩B)

Probability of top wire success P(A ∩ B) = P(A)P(B) = $\frac{16}{25}$
Probability of middle wire success P(A)  = $\frac{2}{3}$
Probability of lowest wire success P(A ∩ B) = $\frac{9}{16}$

Probability of circuit success = P(A U B U C) = P(A) + P(B) + P(C) - P(A ∩ B) - P(A ∩ C) - P(B ∩ C) + P(A ∩ B ∩ C)
                                                                   =$\frac{16}{25} + \frac{2}{3} + \frac{9}{16} - \frac{16}{25}* \frac{2}{3} - \frac{16}{25}*\frac{9}{16} - \frac{2}{3}* \frac{9}{16} + \frac{16}{25}*\frac{2}{3}* \frac{9}{16}$
 
                                                                    = $\frac{379}{400}$ OPTION B

• edited by
0 0 votes

We need to find the probability that $A$ and $B$ are connected, which is denoted as $P(A \text{ and } B \text{ connected})$.

Let:

  • Top Wire be $W_1$

  • Middle Wire be $W_2$

  • Bottom Wire be $W_3$

The terminals $A$ and $B$ will be connected if at least one of the wires is working.

$$P(A \text{ and } B \text{ connected}) = P(W_1 \text{ working} \cup W_2 \text{ working} \cup W_3 \text{ working})$$

Using the complement rule, this can be written as:

$$P(A \text{ and } B \text{ connected}) = 1 - P(\text{none of the wires are working})$$

$$P(A \text{ and } B \text{ connected}) = 1 - P(\overline{W_1} \cap \overline{W_2} \cap \overline{W_3})$$

Since the failure probabilities of the links are independent:

$$P(\overline{W_1} \cap \overline{W_2} \cap \overline{W_3}) = P(\overline{W_1}) \cdot P(\overline{W_2}) \cdot P(\overline{W_3})$$

For Wire 1 ($W_1$):

$$P(\overline{W_1}) = P(\text{point 1 not working} \cup \text{point 2 not working})$$

$$P(\overline{W_1}) = 1 - P(\text{point 1 working} \cap \text{point 2 working})$$

$$P(\overline{W_1}) = 1 - \left(P(\text{point 1 working}) \cdot P(\text{point 2 working})\right)$$

$$P(\overline{W_1}) = 1 - \left[1 - \frac{1}{5}\right] \cdot \left[1 - \frac{1}{5}\right]$$

$$P(\overline{W_1}) = 1 - \frac{4}{5} \cdot \frac{4}{5}$$

$$P(\overline{W_1}) = 1 - \frac{16}{25}$$

$$\therefore P(\overline{W_1}) = \frac{9}{25}$$

For Wire 2 ($W_2$):

$$P(\overline{W_2}) = \frac{1}{3}$$

For Wire 3 ($W_3$):

$$P(\overline{W_3}) = P(\text{point 1 not working} \cup \text{point 2 not working})$$

$$P(\overline{W_3}) = 1 - \left(P(\text{point 1 working}) \cdot P(\text{point 2 working})\right)$$

$$P(\overline{W_3}) = 1 - \left[1 - \frac{1}{4}\right] \cdot \left[1 - \frac{1}{4}\right]$$

$$P(\overline{W_3}) = 1 - \frac{3}{4} \cdot \frac{3}{4}$$

$$P(\overline{W_3}) = 1 - \frac{9}{16}$$

$$\therefore P(\overline{W_3}) = \frac{7}{16}$$

Calculating the final probability:

$$\therefore P(\text{none of the wires working}) = \frac{9}{25} \cdot \frac{1}{3} \cdot \frac{7}{16}$$

$$P(\text{none of the wires working}) = \frac{3}{25} \cdot \frac{7}{16} = \frac{21}{400}$$

Now, substitute this back into our original equation:

$$\therefore P(A \text{ and } B \text{ connected}) = 1 - \frac{21}{400}$$

$$P(A \text{ and } B \text{ connected}) = \frac{379}{400}$$

Therefore, the correct option is B.

Answer:
Position:
Show:

Related questions

23 23 votes
5 answers 5 answers
5.5k
5.5k views
Misbah Ghaya asked Oct 30, 2015
5,486 views
There are $1000$ balls in a bag, of which $900$ are black and $100$ are white. I randomly draw $100$ balls from the bag. What is the probability that the $101$st ball wil...
11 11 votes
6 6 answers
4.4k
4.4k views
Misbah Ghaya asked Oct 30, 2015
4,425 views
A spider is at the bottom of a cliff, and is $n$ inches from the top. Every step it takes brings it one inch closer to the top with probability $1/3$, and one inch away f...
14 14 votes
2 answers 2 answers
3.0k
3.0k views
Misbah Ghaya asked Oct 26, 2015
3,004 views
The probability of throwing six perfect dices and getting six different faces is$1 -\dfrac{ 6!} { 6^{6}}$$\dfrac{6! }{ 6^{6}}$$6^{-6}$$1 - 6^{-6}$None of the above
22 22 votes
11 answers 11 answers
14.4k
14.4k views
Misbah Ghaya asked Oct 25, 2015
14,413 views
Amar and Akbar both tell the truth with probability $\dfrac{3 } {4}$ and lie with probability $\dfrac{1}{4}$. Amar watches a test match and talks to Akbar about the outco...